OpenStudy (anonymous):

will give a medal and be fanned Which operations are not commutative? division and addition multiplication and subtraction division and subtraction addition and subtraction

OpenStudy (anonymous):

Because this is a mathematical question! :P But here, Commutative Operation Any operation for which a+b = b+a. Addition and multiplication are both commutative. Subtraction, division, and composition of functions are not. For example, 5 + 6 = 6 + 5 but 5 – 6 ≠ 6 – 5.

OpenStudy (anonymous):

So the answer would be C. Division and Subtraction are not commutative operations

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

gave you a medal